### v3

# arXiv:2505.13064v3[math.DS] 02 Mar 2026

Abstract
Multi-body mechanical systems have rich internal dynamics, whose solutions can be exploited as efficient control targets. Yet, solutions non-trivially depend on system parameters, obscuring feasible properties for use as target trajectories. For periodic regulation tasks in robotics applications, we investigate properties of nonlinear normal modes (NNMs) collected in Lyapunov subcenter manifolds (LSMs) of conservative mechanical systems. Using a time-symmetry of conservative mechanical systems (CMs), we show that mild non-resonance conditions guarantee LSMs to be Eigenmanifolds, in which NNMs are guaranteed to oscillate between two points of zero velocity. We also prove the existence of a unique generator, which is a connected, 1D manifold that collects these points of zero velocity for a given Eigenmanifold. Furthermore, we show that an additional spatial symmetry provides LSMs with yet stronger properties of Rosenberg manifolds. Here all brake trajectories pass through a unique equilibrium configuration, which can be favorable for control applications. These theoretical results are numerically confirmed on two mechanical systems: a double pendulum and a 5-link pendulum.
keywords:
Nonlinear Dynamics, Multi-body mechanics, Eigenmanifolds, Symmetries, Differential Geometry
thanks: This work is supported by the Advanced Grant M-Runner (Grant Agreement No. 835284) by the European Research Council (ERC).
This work is supported by the Advanced Grant PortWings (Grant Agreement No. 787675) by the ERC.
1 Introduction
Eigenmanifolds
Rosenberg manifolds
$\mathcal{Q}$
$\bar{x}$
$\left(x,0\right)$
$\left(x(t),P(t)\right)$
Generator
Generator
$\left(\varphi(x),0\right)$
Linearized Mode
$\mathcal{Q}$
$\bar{x}$
$\left(x,0\right)$
$\left(x(t),P(t)\right)$
Linearized Mode
Generator
Generator
$\mathcal{Q}$
$\bar{x}$
$\left(x(t),P(t)\right)$
Family of periodic orbits
Family of eigenmodes
$\mathcal{Q}$
$\bar{x}$
$\left(x(t),P(t)\right)$
General periodic orbit
$\mathcal{Q}$
$\bar{x}$
$\left(x,0\right)$
$\left(x(t),P(t)\right)$
Geometric eigenmode
Family of Rosenberg modes
$\mathcal{Q}$
$\bar{x}$
$\left(x,0\right)$
$\left(\varphi(x),0\right)$
$\left(x(t),P(t)\right)$
Geometric Rosenberg mode
Thms.
Thm. 12
4, 7
a)
b)
c)
d)
e)
f)
Lyapunov sub-
center manifolds
Figure 1: Summary of the article: for conservative mechanical systems with configuration $x\in\mathcal{Q}$, momentum $P\in T^{*}_{x}\mathcal{Q}$, and projecting all data to $\mathcal{Q}$ for ease of visualization. Lyapunov subcenter manifolds (panels a and d) are families of general periodic oscillations $(x(t),P(t))$ springing from an equilibrium $(\bar{x},0)$. We prove conditions for LSMs to have stronger properties: Theorems 4 and 7 make it highly common for LSMs to become weak Eigenmanifolds (panels b and e) that collect periodic brake trajectories (oscillating between brake points $(x,0)$), and Eigenmanifolds that collect periodic brake trajectories whose configuration trajectory does not self-intersect (called geometric eigenmodes). Theorem 12 shows that yet stronger conditions turn LSMs into (weak) Rosenberg manifolds (panels c and f), where all modal configurations pass through $\bar{x}$. In both (weak) Eigenmanifolds and (weak) Rosenberg manifolds, brake points are collected on a connected 1D submanifold that we call the generator.
Practical advances in robotics are pushing the need for a deeper understanding of the nonlinear dynamic behavior of multibody mechanical systems [1, 2, 3]. For example, a key challenge on the frontier of robotics lies in enabling robots to move efficiently while performing periodic motions (such as repetitive industrial operations and locomotion), thus avoiding motor torque limits and mitigating limited battery capacity [4, 5, 6, 7, 8].
A growing body of literature addresses this challenge by exciting internal robot dynamics for low-torque control implementations, with modal analysis serving as the basis for these works [9, 10, 11, 12, 13].Given a system with state $\zeta$, control-input $u$ representing a motor torque, current or voltage and assuming control-affine dynamics
$\dot{\zeta}=f(\zeta)+g(\zeta)u\,.$
Then the most energy-efficient control-input is simply $u=0$: when tracking target trajectories that are natural motions of $f(\zeta)$, this corresponds to the steady-state control-input, forming the rationale behind exciting internal robot dynamics. For periodic regulation tasks, (periodic) natural motions from a dissipation-free model serve as target trajectories for the real system, where stabilizing controllers compensate for small dissipative effects but not the much larger inertial terms [14, 15].
To use natural motions as control-targets in practice, optimization of robots and their internal dynamics becomes a core component [16]. In this optimization of design parameters it is useful to know what properties can or cannot be expected from the natural motions of the nonlinear and often chaotic open-loop dynamics [17].
However, it is challenging to derive such knowledge from direct application of classic modal theory from nonlinear mechanics [18, 19, 20]: robots are usually multibody systems whose configuration is modeled to evolve on Riemannian manifolds with curvature [21], which fall outside standard theory, leading instead to postulated geometric definitions [22].
In this article we go back to the foundations of modal theory in nonlinear dynamics, starting from Lyapunov subcenter manifold (LSM) theory to describe the families of oscillations that can be expected in multibody dynamical systems, and to rigorously derive robust, geometric properties.
Given a dynamical system $\dot{\zeta}=f(\zeta)$ with a conserved quantity $H(\zeta)$ (i.e., $\dot{H}=0$), LSMs are two-dimensional submanifolds of the state-space that originate at equilibria with imaginary eigenvalues and collect periodic trajectories [23]. LSMs have been heavily investigated over the past decades [24, 25, 26] and have found important applications to disparate scenarios [27, 28, 29]. LSMs are well-studied: there are results on existence, uniqueness, differentiability [30, 31], and persistence under dissipative disturbances [32] in the form of spectral submanifolds.
The theory of LSMs directly applies to conservative mechanical systems (CMs), i.e., including conservative multi-body mechanical systems, for which the Hamiltonian
$H(x,P)=\frac{1}{2}M(x)^{-1}(P,P)+V(x)$ (1)
is the conserved quantity – where we used $x\in\mathcal{Q},P\in T^{*}_{x}\mathcal{Q}$ to denote abstract configuration and momentum variables, denote as $M(x)$ the symmetric, positive-definite inertia tensor, and $V:\mathcal{Q}\rightarrow\mathbb{R}$ the potential energy. It is well-known that unique LSMs exist around equilibria of CMs (i.e., minima of $V$) when the eigenvalues of the linearized system full-fill a non-resonance condition. However, high-level properties of the periodic trajectories collected in LSMs of CMs are apriori undetermined in this framework, see Figure 1 a & d.
Our contribution utilizes the time-symmetry inherent to CMs to shows that unique LSMs collect families of periodic brake trajectories, which are trajectories whose configuration oscillates back and forth between points with zero momentum. Uniqueness of LSMs is a robust feature over large ranges of parameters, and we argue that families of brake trajectories are a feasible target trajectories for e.g., optimization of pick-and-place tasks. In line with related literature that defined families of periodic brake trajectories without deriving them from first principles [22, Sec. 7], we refer to these families as Eigenmanifolds, see Figure 1 b & e.
Furthermore, we show that an additional spatial symmetry gives LSMs the stronger properties of Rosenberg manifolds [22], see Figure 1 c & f. In these structures all oscillation pass through the equilibrium configuration, which is a desirable control property since it makes it easy to switch between modes by impulsive control actions [33]. We derive conditions on the inertia-tensor and potential energy function of CMs, for the existing LSMs to be Rosenberg manifolds. In local coordinates $q,p$ with equilibrium at $q=0$, inertia matrix $M(q)\in\mathbb{R}^{n\times n}$ and potential energy $V(q)\in\mathbb{R}$, these conditions can be enforced as M(q)=M(q)M(q)=M(-q) and V(q)=V(q)V(q)=V(-q). This makes it possible to constrain Eigenmanifolds into Rosenberg manifolds at a design stage, enabling stricter properties of feasible target trajectories.
In summary, our contributions are:
  1. 1.
    Theorem 5 properties of LSMs induced by symmetries of dynamic systems, and the propagation of symmetries to individual trajectories in LSMs
  2. 2.
    Various results on LSMs in CMs collected in Theorem 7, stating that unique LSMs collect periodic brake trajectories, and admit generators, which lends itself to a convenient parametrization of LSMs.
  3. 3.
    Theorem 12 that shows that conditions M(q)=M(q)M(q)=M(-q) and V(q)=V(q)V(q)=V(-q) let (weak) Eigenmanifolds in CMs exhibit the stronger properties of (weak) Rosenberg manifolds.
The rest of the paper is organized as follows. Sec. 2 introduces background theory on CMs, LSMs and discrete symmetries. Sec. 3 presents conditions on $M(q)$ and $V(q)$ that determine existence and uniqueness of LSMs in CMs, and shows conditions under which self-symmetric LSMs consist of self-symmetric periodic motions. Sec. 4 derives the main theoretical contributions of this article, properties of self-symmetric LSMs and self-symmetric periodic motions in CMs. For periodic orbits, this motivates our definitions of geometric eigenmodes and geometric Rosenberg modes. For LSMs, we prove that unique LSMs of CMs are collections of geometric eigenmodes and admit a generator, by which we justify the definition of Eigenmanifolds. We derive a sufficient condition for Eigenmanifolds to fulfill the stronger property that all trajectories pass through the equilibrium configuration, using this to connect to nonlinear mechanics definitions of Rosenberg manifolds.
Sec. 5 verifies the results in numerical examples of CMs, highlighting that the existence of Eigenmanifolds and Rosenberg manifolds follows from rules implemented at a design stage.
We end with a conclusion in Sec. 6. Proofs are given in Appendix A, and Appendix B (only in the supplementary material [34]) contains background on choice of coordinates.
1.1 Related literature
Exciting and exploiting efficient nonlinear oscillations is an open challenge in nonlinear control theory [35, 36, 37], with important ramifications in robotics’ locomotion and execution of industrial operations [38, 39, 40, 41, 42], and biomechanics [43, 44, 45]. To this end, Eigenmanifolds have been recently introduced (see [22, Sec. 7]) as an extension of modal analysis to multi-body systems, in a way that naturally lends itself to control applications: design of mechanical systems [46] and shaping of Eigenmanifolds via the systems potential energy [47, 16] has been explored. Eigenmanifolds of interest are stabilized in simulation [48, 49], on real systems [50], can be stabilized in an energy-efficient manner [15] and allow the swing-up of weakly actuated systems to high energies [14]. Eigenmanifolds have also been applied in classifying the accuracy of reduced order models in soft robotics [51, 52].
Eigenmanifold theory has its precursors in the study of nonlinear normal modes [53, 20, 54] for nonlinear mechanical systems with constant inertia tensors. Rosenberg modes are well-studied for conservative mechanical systems [18, 19, 55], and their symmetry properties are known by the condition that V(q)=V(q)V(q)=V(-q). Strict normal modes are a special case of Eigenmanifolds that exist under compatibility of the potential gradient with the geodesic flow [56].
Within the nonlinear dynamics literature, LSMs are long studied [23, 30, 31] and their persistence in the limit of small dissipation was shown [32], but this has not been connected to insights about Eigenmanifolds in multi-body dynamics. LSMs have also been applied to the model order reduction of partial differential equations [57].
A general treatment of discrete symmetries in dynamical systems is given by Michael Field [58]. Large classes of symmetries were also studied in the context of periodic orbits and nonlinear normal modes [59, 60]. However, these works largely treat symmetries on vector and Banach spaces, and do not directly apply to Eigenmanifold theory. An excellent overview to time-reversal symmetries in Hamiltonian systems is given by Lamb & Roberts [61], covering many properties of direct relevance to Eigenmanifold theory, such as the existence of time-symmetric LSMs. Alomair & Montaldi [62] investigate the interplay of a time-reversal and a spatial symmetry of Hamiltonian systems in the context of periodic orbits, proving the existence of a discrete number of doubly symmetric periodic orbits around symmetric equilibria. However, both of [61, 62] leave properties of such symmetric orbits and their families implicit.
2 Preliminaries and problem statement
We briefly introduce the Hamiltonian description of conservative mechanical systems, Lyapunov subcenter manifolds, and discrete symmetries. We conclude with a formal problem statement.
2.1 Notation
Table 1: Table of Symbols and Notation
Symbol Description Defined in
$\mathcal{Q}$ configuration manifold Sec. 2.2
$T^{*}\mathcal{Q}$ cotangent bundle Sec. 2.2
$\mathcal{M}\subset T^{*}\mathcal{Q}$ LSM, Eigenmanifold or Rosenberg manifold Secs. 2.2, 2.3
$\mathcal{R}\subset\mathcal{M}$ generator of an Eigenmanifold or Rosenberg manifold Sec. 2.2
$x\in\mathcal{Q},\,P\in T^{*}_{x}\mathcal{Q},\,\zeta=(x,P)\in T^{*}\mathcal{Q}$ abstract configuration and momentum variables Sec. 2.2
$q,p\in\mathbb{R}^{n},\,z=(q,p)\in\mathbb{R}^{2n}$ configuration and momentum in a canonical chart Sec. 2.2
$H:T^{*}\mathcal{Q}\rightarrow\mathbb{R}$ Hamiltonian Sec. 1
$X_{H}\in\mathfrak{X}(T^{*}\mathcal{Q})$ abstract Hamiltonian vector field Sec. 2.2
$\lambda_{0}\in\mathbb{C}$ Eigenvalue of linearized system Sec. 2.3
$E_{0}\in T_{(\bar{x},0)}T^{*}\mathcal{Q}$ Eigenspace of linearized system Sec. 2.3
$\phi_{\gamma}:[0,1]\rightarrow\mathcal{Q}$ immersion or immersion of geometric Eigenmode $\zeta$ Sec. 2.2
$\tau\in\mathbb{R}$ time-scaling Sec. 2.4
$\sigma:\mathcal{M}\rightarrow\mathcal{M}$ symmetry map Sec. 2.4
$S:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ coordinate version of symmetry map Sec. 2.4
$\text{Fix}(\sigma)$ fixed set of a symmetry Sec. 2.4
$\Psi^{t}_{f}:\mathcal{M}\rightarrow\mathcal{M}$ Flow of $f\in\mathfrak{X}(\mathcal{M})$ Sec. 2.4
Smooth manifolds are denoted $\mathcal{M},\mathcal{N},\mathcal{R},\mathcal{Q}$. At a point $x\in\mathcal{M}$, $T_{x}\mathcal{M}$ denotes the tangent space of $\mathcal{M}$ at $x$, and $T^{*}_{x}\mathcal{M}$ denotes the respective cotangent space. The tangent bundle over $\mathcal{M}$ is $T\mathcal{M}$, and the cotangent bundle is $T^{*}\mathcal{M}$. The set of vector fields over $\mathcal{M}$ is $\mathfrak{X}(\mathcal{M})$. For a smooth map $\sigma:\mathcal{M}\rightarrow\mathcal{N}$ between smooth manifolds $\mathcal{M},\mathcal{N}$ we write $\sigma\in C^{\infty}(\mathcal{M},\mathcal{N})$. The pushforward of $\sigma$ is $\sigma_{*}:T_{x}\mathcal{M}\rightarrow T_{\sigma(x)}\mathcal{N}$, and the pullback is $\sigma^{*}:T^{*}_{\sigma(x)}\mathcal{N}\rightarrow T^{*}_{x}\mathcal{M}$. The image of $\sigma$ is denoted $\sigma(\mathcal{M}):=\{\sigma(x)\;|\;x\in\mathcal{M}\}\subseteq\mathcal{N}$.
Additional symbols introduced in the paper are summarized in Table 1.
2.2 Conservative mechanical systems
We describe CMs with configuration space $\mathcal{Q}$ by their Hamiltonian dynamics on the cotangent bundle $T^{*}\mathcal{Q}$, where the Hamiltonian $H$ in Equation (1) canonically determines the Hamiltonian vector field $X_{H}\in\mathfrak{X}(T^{*}\mathcal{Q})$ (for details on the construction, see e.g., [21]). Let $\zeta:\mathbb{R}\rightarrow T^{*}\mathcal{Q}$ denote a solution to the abstract Hamiltonian dynamics
$\dot{\zeta}=X_{H}(\zeta)\,,\;\zeta(0)=\zeta\,.$ (2)
We refer to (2) as a CM when $H$ is given by (1). We denote the corresponding flow of $X_{H}$ as
$\Psi_{X_{H}}^{t}:T^{*}\mathcal{Q}\rightarrow T^{*}\mathcal{Q}\,,\;\zeta(t)=\Psi_{X_{H}}^{t}(\zeta_{0})\,$, (3)
where $\Psi_{X_{H}}^{t}$ is defined such that $\zeta(t)=\Psi_{X_{H}}^{t}(\zeta_{0})$ solves (2). We assume that $X_{H}$ is complete. The set of points on $\zeta(t)$ is
$\zeta(\mathbb{R}):=\{\zeta(t)\;|\;t\in\mathbb{R}\}\,.$ (4)
In a canonical chart (see Appendix B.1 in the supplementary material [34]) with coordinates $q,p\in\mathbb{R}^{n}$, the Hamiltonian dynamics of a CM follow from (1), (2) as:
$\dot{q}$ $=\frac{\partial H}{\partial p}(q,p)=M(q)^{-1}p\,$, (5)
$\dot{p}$ $=-\frac{\partial H}{\partial q}(q,p)=-\frac{\partial}{\partial q}\left(\frac{1}{2}p^{\top}M(q)^{-1}p\right)-\frac{\partial V}{\partial q}(q)\,.$ (6)
We are interested in families of periodic oscillations springing from an isolated, stable equilibrium of (5), (6) (i.e., a point $(\bar{q},\bar{p})$ with $\bar{p}=0$, $\frac{\partial V}{\partial q}(\bar{q})=0$, and $\frac{\partial^{2}V}{\partial q\partial q}(\bar{q})>0$). In a coordinate-free setting we denote this equilibrium as $\bar{\zeta}=(\bar{x},0)$, and in a chart-based setting we choose coordinates such that the equilibrium is $\bar{z}=(\bar{q},0)=(0,0)$.
2.3 Lyapunov subcenter manifolds
We briefly introduce Lyapunov subcenter manifolds (LSMs), focusing on applications to systems of the type (5), (6). For a complete treatment, refer to [32, 31] and references therein.
Given any dynamical system in local coordinates $z\in\mathbb{R}^{2n}$
$\dot{z}=Az+B(z)\,$, (7)
where $A\in\mathbb{R}^{2n\times 2n}$, $B:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$ is analytic and $B(0)=0$, $\frac{\partial B}{\partial z}(0)=0$.
Denote as $(\lambda_{1},\ldots,\lambda_{2n})$ the eigenvalues of $A$, and define the linear eigenspace associated with $\lambda_{k}$
$E_{k}:=\ker(\lambda_{k}I-A)\subset\mathbb{C}^{2n}\,.$ (8)
We repeat [32, Ass. 2.1 & Thm. 2.5] on the existence and uniqueness of LSMs:
Theorem 1 (Lyapunov subcenter manifolds).
If it holds that
  1. (i)
    A is diagonalizable.
  2. (ii)
    A has a pair of conjugate eigenvalues $\pm i\omega_{0}$ with zero real part.
  3. (iii)
    The remaining 2n22n-2 eigenvalues $\lambda_{k}$ of A, with $1\leq k\leq 2n-2$, are non-resonant with the eigenvalues $\pm i\omega_{0}$, i.e., it holds for all $k$ that
    $\frac{\lambda_{k}}{i\omega_{0}}\notin\mathbb{Z}\,.$ (9)
  4. (iv)
    The system (7) has a conserved quantity $H(z(t))$ along integral curves $z(t)$, i.e.,
    $\frac{d}{dt}H(z(t))=0\,.$ (10)
    Further, $H(0)=0$, $(\frac{\partial}{\partial z}H)(0)=0$, and
    $Y^{\top}\left(\left(\frac{\partial^{2}}{\partial z\partial z^{T}}H\right)(0)\right)Y>0\;,\forall Y\in E_{0},$
    with $E_{0}$ the eigenspace associated with $\pm i\omega_{0}$.
Then there exists, locally, a unique 2D submanifold $\mathcal{M}\subseteq\mathbb{R}^{2n}$ of periodic trajectories tangent to the eigenspace $E_{0}$ of $\pm i\omega_{0}$.
This 2D submanifold $\mathcal{M}$ is called the Lyapunov subcenter manifold associated with the eigenspace $E_{0}$ at $z=0$. Apart from periodicity, there are no conditions on trajectories collected within an LSM (cf. Figure 1a). When the non-resonance condition is not fulfilled, Lyapunov subcenter manifolds can still be found, but they are not unique [30].
It follows by repeated application of Theorem 1:
Corollary 2.
If $d$ of the $n$ pairs of eigenvalues $\pm i\omega_{k}$ of $A$, $1\leq k\leq n$, are mutually non-resonant, and non-resonant with the remaining eigenvalues, then there exist, locally, at least $d$ unique families of periodic orbits, with a unique family being tangent to a given eigenspace $E_{k}$ of $A$.
Note that Corollary 2 provides a lower bound since the conditions of Theorem 1 are sufficient, but not necessary.
2.4 Symmetries of CMs
We briefly introduce discrete symmetries of Hamiltonian dynamics, as well as a well-known time symmetry of CMs. For a general treatment of symmetries of Hamiltonian systems see e.g., [61, 62].
Definition 3 (Discrete Symmetry).
A symmetry $(\sigma,\tau)$ of the Hamiltonian dynamics $X_{H}\in\mathfrak{X}(T^{*}\mathcal{Q})$ consists of a diffeomorphism $\sigma:T^{*}\mathcal{Q}\rightarrow T^{*}\mathcal{Q}$ and a scalar $\tau\in\{-1,1\}$ such that
$X_{H}\circ\sigma=\tau\sigma_{*}X_{H}\,.$ (11)
It is a well-known result that symmetries map solutions to solutions:
$\Psi^{t}_{X_{H}}\big(\sigma(\zeta)\big)$ $=\sigma\big(\Psi^{\tau t}_{X_{H}}(\zeta)\big)\,$, (12)
i.e., for any solution $\zeta(t)$ of the dynamics $X_{H}$, also $\sigma(\zeta(\tau t))$ is a solution.
We define the fixed orbit subset $\text{Fix}(\sigma)\subset T^{*}\mathcal{Q}$ as the union of orbits of $X_{H}$} that are fixed under a symmetry $(\sigma,\tau)$
$\text{Fix}(\sigma):=\{\zeta\in\Psi^{\mathbb{R}}_{X_{H}}(\zeta_{0})\;|\;\Psi^{\mathbb{R}}_{X_{H}}(\zeta_{0})=\sigma\Psi^{\mathbb{R}}_{X_{H}}(\zeta_{0})\}\,.$ (13)
In a local chart $(U,Z)$ with $\zeta,\sigma(\zeta)\in U\subset T^{*}\mathcal{Q}$ and $Z:U\subset T^{*}\mathcal{Q}\rightarrow\mathbb{R}^{2n}$, the vector field $X_{H}$ is represented by components $f^{i}(z)\in\mathbb{R}$, $i\in\{1,\cdots,n\}$, and the symmetry as a map $S:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$ with $S=Z\circ\sigma\circ Z^{-1}$. Then condition (11) reads
$f^{i}(S(z))=\tau\frac{\partial S^{i}}{\partial z^{j}}f^{j}(z)\,.$ (14)
If this holds, then for any solution $z(t)$ of the dynamics $\dot{z}^{i}=f^{i}(z)$, also $S\big(z(\tau t)\big)$ is a solution.
The equations of motion of a CM with Hamiltonian dynamics (2) are subject to the time-reversal symmetry $(x,P,t)\rightarrow(x,-P,-t)$. As shown by [61], this follows more generally for any Hamiltonian system when H(x,P)=H(x,P)H(x,P)=H(x,-P). Following Definition 3, we write this as $({\sigma}_{1},\tau_{1})$, with
${\sigma}_{1}\big((x,P)\big)=(x,-P)\,,\quad\tau_{1}=-1\,.$ (15)
As a consequence of this symmetry, for any solution $\zeta(t)=(x(t),P(t))$ of (2) with Hamiltonian (1), also $\sigma_{1}(\zeta(\tau_{1}t))=(x(-t),-P(-t))$ is a solution of (2). Note that $\sigma_{1}$ leaves any local subspace $E\subset T_{\zeta}T^{*}\mathcal{Q}$ invariant. We will later show that this means that $\sigma_{1}$ maps any LSM into itself if the LSM is uniquely associated with a given eigenspace, which we then show will induce additional properties.
2.5 Problem statement
In this article, we aim to answer the following questions:
  1. 1.
    How do the conditions for existence and uniqueness of LSMs translate to conditions on the inertia tensor $M(q)$ and the potential energy $V(q)$?
  2. 2.
    What high-level properties of periodic trajectories in LSMs follow from the symmetry (15)?
  3. 3.
    What other symmetries, corresponding to conditions on $M(q)$ and $V(q)$, result in desirable properties of LSMs?
3 Lyapunov subcenter manifolds in conservative mechanical systems
3.1 Conditions on inertia tensor and potential energy
We present Theorem 4, containing conditions on the inertia tensor $M(q)$ and the potential energy $V(q)$ that guarantee the existence and uniqueness LSMs at an equilibrium $(\bar{x},0)$. This is a minor contribution through its geometric (i.e., chart-free) formalism.
Theorem 4 (LSMs in mechanical systems).
Given the Hamiltonian dynamics (2) on $T^{*}\mathcal{Q}$ with analytic Hamiltonian (1). If there is a point $\bar{x}\in\mathcal{Q}$ where it holds that
  1. (i)
    $\text{d}V(\bar{x})=0$ and $\text{Hess}(V)(\bar{x})>0$
  2. (ii)
    The $(0,2)$-tensor $M(\bar{x})$ is positive definite and symmetric
  3. (iii)
    The $(1,1)$-tensor $M(\bar{x})^{-1}\text{Hess}(V)(\bar{x})$ is diagonalizable
  4. (iv)
    The eigenvalue $\lambda_{0}^{2}=-\omega_{0}^{2}$ of $M(\bar{x})^{-1}\text{Hess}(V)(\bar{x})$ is non-resonant with the remaining eigenvalues $\lambda_{k}$, $1\leq k\leq n-1$, which means that
    $\frac{\lambda_{k}^{2}}{\lambda_{0}^{2}}\notin\mathbb{Z}\,.$ (16)
Then there exists, locally, a unique 2D submanifold $\mathcal{M}\subseteq T^{*}\mathcal{Q}$ of periodic trajectories which is tangent to the eigenspace $E_{0}=D_{0}\oplus M(\bar{x})D_{0}\subset T_{(\bar{\zeta})}T^{*}\mathcal{Q}$, where $D_{0}=\ker(\omega_{0}^{2}I-M(\bar{x})^{-1}\text{Hess}(V)(\bar{x}))$.
{pf}
The proof relies on showing that the conditions of Theorem 4 are equivalent to those of Theorem 1.
First, express the Hamiltonian dynamics of a conservative mechanical system (2), (6) in canonical coordinates $(q,p)$ such that the equilibrium $(\bar{x},0)$ is mapped to $q=0$, $p=0$. Further write $z=(q,p)\in\mathbb{R}^{2n}$. The dynamics then take the form
$\dot{z}$ $:=f(z)=\begin{pmatrix}M(q)^{-1}p\\ -\frac{\partial}{\partial q}(p^{\top}M(q)^{-1}p)-\frac{\partial V}{\partial q}(q)\end{pmatrix}\,$, (17)
which can be rewritten as
$\dot{z}=Az+B(z)\,$, (18)
with
$A$ $=\frac{\partial f}{\partial z}\big((0,0)\big)=\begin{bmatrix}0&M(0)^{-1}\\ -\frac{\partial^{2}V}{\partial q\partial q}(0)&0\end{bmatrix}\,$, (19)
$B(z)$ $=f(z)-Az\,.$ (20)
The dynamics are of the required form (7) for Theorem 1, since $B(0)=0$, and $\frac{\partial B}{\partial z}(0)=0$. $B(z)$ is also analytic, given that $M(q),V(q)$ are analytic. This does not rely on the choice of a particular chart: the components of $A$ correspond to the components of a $(1,1)$-tensor on $T^{*}\mathcal{Q}$, given that $\text{d}V(\bar{x})=0$, which follows from $(0,0)$ being an equilibrium of the dynamics (5), (6).
The eigenvalues $\lambda$ of $A=\begin{bmatrix}0&B\\ C&0\end{bmatrix}$ are the solution to
$\det(\lambda I_{2n}-A)=\det(\lambda^{2}I_{n}-BC)=0\,$, (21)
such that $A$ is diagonalizable since the $(1,1)$-tensor $BC=-M(0)^{-1}\frac{\partial^{2}V}{\partial q\partial q}(0)$ is diagonalizable. Thus, condition $(i)$ of Theorem 1 is fulfilled.
Further, $\lambda^{2}=-\omega^{2}$ for some positive $\omega\in\mathbb{R}_{+}$, as $M(0)^{-1}\frac{\partial^{2}V}{\partial q\partial q}(0)$ is a positive definite matrix collecting the components of the positive definite tensor $M(\bar{x})^{-1}\text{Hess}(V)(\bar{x})$, and eigenvalues $\omega_{j}^{2}$ of $M(0)^{-1}\frac{\partial^{2}V}{\partial q\partial q}(0)$ correspond to pairs of eigenvalues
$\lambda_{j,\pm}=\pm i\omega_{j}$ (22)
of $A$. Thus, condition $(ii)$ of Theorem 1 is fulfilled.
The eigenvalues $\lambda_{0,\pm}$ are non-resonant with the remaining 2n22n-2 eigenvalues $\lambda_{k,\pm}$ of $A$, since
$\frac{\lambda_{k}^{2}}{\lambda^{2}_{0,\pm}}\notin\mathbb{Z}\Rightarrow\frac{\lambda_{k}}{\lambda_{0,\pm}}\notin\mathbb{Z}\,$, (23)
where $\frac{\lambda_{k}^{2}}{\lambda^{2}_{0,\pm}}\notin\mathbb{Z}$ is guaranteed by assumption. Thus, also condition $(iii)$ of Theorem 1 is fulfilled.
Finally, a conserved quantity $H\big(z(t)\big)$ fulfilling condition $(iv)$ of Theorem 1 is given by the Hamiltonian $H\big(z(t)\big)=\frac{1}{2}p^{\top}M(q)^{-1}p+V(q)$ .
The conditions of Theorem 4 are chart-invariant: the equilibrium is guaranteed to be of the form $(\bar{x},0)$ in any canonical chart, $M(q)$ is positive definite and symmetric in any chart, and eigenvalues of the component matrix $M(0)^{-1}\frac{\partial^{2}V}{\partial q\partial q}(0)$ of the $(1,1)$-tensor $M(\bar{x})^{-1}\text{Hess}(V)(\bar{x})$ are chart-invariant.
3.2 Propagation of symmetries
Theorem 5 presented hereafter, is a novel result that is essential to derive results in the remainder of this section. Given a symmetry $(\sigma,\tau)$ of $X_{H}$, it presents conditions for an LSM $\mathcal{M}$ and trajectories contained within it to be fixed sets of the symmetry.
Theorem 5 (Symmetric LSMs).
[…]
$\sigma((\bar{x},0))=(\bar{x},0)\,$, (24)
and
$E_{0}=\sigma_{*}E_{0}\,$, (25)
then it also holds that
$\mathcal{M}=\sigma\big(\mathcal{M}\big)\,$, (26)
and any evolution $\zeta(\mathbb{R})\in\mathcal{M}$ satisfies
$\zeta(\mathbb{R})=\sigma(\zeta(\mathbb{R}))\,.$ (27)
{pf}
Begin by showing that $\mathcal{M}=\sigma(\mathcal{M})$. The presence of the symmetry $\sigma$ implies that both $\mathcal{M}$ and $\mathcal{M}^{\prime}:=\sigma(\mathcal{M})$ are LSMs, and condition $\sigma(0)=0$ implies that $0\in\mathcal{M}$ and $0\in\mathcal{M}^{\prime}$, i.e., both are LSMs about the equilibrium $z=0$. Further, ${\sigma}_{*}E_{0}=E_{0}$ implies that both $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are tangent to the eigenspace $E_{0}$. Theorem 4 holds, such that $\mathcal{M}$ is the unique LSM tangent to $E_{0}$, and hence it must be that $\mathcal{M}=\mathcal{M}^{\prime}$.
To show that also $\zeta(\mathbb{R})=\sigma(\zeta(\mathbb{R}))$, introduce coordinates $(U,X)$ on $\mathcal{M}$, i.e., $U\subset\mathcal{M}$ is a neighborhood of $z=0$ and $X:\mathcal{M}\rightarrow\mathbb{R}^{2}$. Further, choose ”polar” coordinates $X$ and a time scaling $\tau:\mathcal{M}\rightarrow\mathbb{R}_{+}$ such that $X(\zeta(t))=(r,t)$ with $r$ constant for a given oscillation $\zeta\in\mathcal{M}$. Then $\sigma:\mathcal{M}\rightarrow\mathcal{M}$ translates into coordinates as
$S(r,t)=(\alpha(r),t+c)\,$, (28)
Since it must map solutions to solutions. The condition
$\sigma\circ\sigma=\text{id}_{\mathcal{M}}\,,\quad\sigma(0)=0$ (29)
translates into coordinates as
$S(S(r,\theta))=(r,\theta)\,,\quad\Phi(0,\theta)=0\,$, (30)
or in terms of $\alpha$
$\alpha(\alpha(r))=r\,,\quad\alpha(0)=0\,$, (31)
With the additional constraint that $\alpha(r)\geq 0$, this has the unique continuous solution $\alpha(r)=r$. Thus
S(r,t)=(r,t+c),S(r,t)=(r,t+c)\,, (32)
necessarily maps solutions into themselves, i.e., $\sigma(\zeta(\mathbb{R}))=\zeta(\mathbb{R})$, completing the proof.
Remark 6.
In a local chart, condition (25) requires that $E_{0}$ is an eigenspace of the Jacobian $\frac{\partial S}{\partial z}_{|z=0}$.
4 Properties of Lyapunov subcenter manifolds induced by symmetry
This section presents our main results, combining Theorems 4, 5 with the symmetry (15) to derive properties of self-symmetric LSMs in CMs. We further present a general class of spatial symmetries that are equivalent to conditions on $M(q),V(q)$ and that lead to desirable control-theoretic properties of LSMs in CMs.
4.1 Time-symmetric Lyapunov subcenter manifolds
We present the properties of time-symmetric Lyapunov subcenter manifold in a theorem:
Theorem 7 (Eigenmanifolds).
Given the conditions of Theorem 4, and with $\pi(x,p)=x$ the projection of $T^{*}\mathcal{Q}$ to $\mathcal{Q}$. Then the unique LSM $\mathcal{M}\subset T^{*}\mathcal{Q}$ associated with the eigenspace $E_{0}\subset T_{(\bar{q},0)}T^{*}\mathcal{Q}$ satisfies the following properties:
  1. 1.
    Immersion: For any trajectory $\zeta(t)\in\mathcal{M}$ there is an immersion [^2] The interval $[0,1]$ is a manifold with boundary, for which the notion of a immersion is not immediately defined. An immersion $\sigma:\mathcal{P}\rightarrow\mathcal{N}$ for a manifold $\mathcal{P}$ with boundary $\partial\mathcal{P}$ is defined such that $\sigma_{|\partial\mathcal{P}}$ is an immersion of $\partial\mathcal{P}$. For $\mathcal{P}=[0,1]$, the boundary is $\partial\mathcal{P}=\{0,1\}$. Since this boundary is 0-dimensional it has no tangent-space, and we define that any map with injective differential on $(0,1)$ classifies as a immersion for the interval $[0,1]$. $\phi_{\zeta}:[0,1]\rightarrow\mathcal{Q}$ such that
    $\phi_{\zeta}([0,1])=\pi(\zeta(\mathbb{R}))$ (33)
  2. 2.
    Embedding: For a neighborhood $\mathcal{M}^{\prime}\subset\mathcal{M}$ around the equilibrium, $\phi_{\zeta}:[0,1]\rightarrow\mathcal{Q}$ is an embedding. When $\dim\mathcal{Q}=2$, then $\mathcal{M}^{\prime}=\mathcal{M}$.
  3. 3.
    Brake points: A trajectory $\zeta(t)\in\mathcal{M}$ encounters two points with $p(t)=0$ over a period of oscillation $T_{\zeta}$, at the configurations $\phi_{\zeta}(0)$ and $\phi_{\zeta}(1)$, and they are a time $T_{\zeta}/2$ apart.
  4. 4.
    Generator: The configurations $\phi_{\zeta}(0)$ and $\phi_{\zeta}(1)$ of the $\zeta\in\mathcal{M}$ lie on a connected, 1D-submanifold $\mathcal{R}\subset\mathcal{M}$ that passes through the equilibrium.
{pf}
See Appendix A.1.
The existence of an immersion $\phi_{\zeta}:[0,1]\rightarrow\mathcal{Q}$ satisfying (33) means that configurations $x(t)$ evolve on a 1D line that may self-intersect. When $\phi_{\zeta}:[0,1]\rightarrow\mathcal{Q}$ is an embedding, then the evolution of the configuration $x(t)$ does not self-intersect.
The existence of points with $P(t)=0$ has immediate implications for application, suggesting that motions collected on LSMs of CMs are interesting for e.g., pick-and-place-like tasks where a manipulator comes to a full stop at two points, but not immediately apply to locomotion, where the system is in continuous motion. This property also allows to reuse insights from numerical continuation methods of nonlinear normal modes [20] which search for new periodic motions only adjusting the initial configuration $x_{0}$ and period $T_{\zeta}$, but not the initial momentum $P_{0}$. This directly generalizes to the numerical continuation of unique LSMs in any type of CM. Finally, the fact that configurations $\phi_{\zeta}(0)$ and $\phi_{\zeta}(1)$ lie on a connected, 1D-submanifold $\mathcal{R}\subset\mathcal{M}$ is convenient for parameterizing motions on time-symmetric LSMs in terms of these 1D-submanifolds.
We make a number of definitions inspired by Theorem 7. The definitions closely reflect standard definitions [63], [20], recent definitions in [22, Sec. 7]. Compared to [22, Sec. 7], they follow from first principles.
Definition 8 (Geometric Eigenmodes).
A geometric eigenmode is a $T$-periodic oscillation $\zeta:\mathbb{R}\rightarrow T^{*}\mathcal{Q}$, such that $\pi(\zeta(\mathbb{R}))\cong[0,1]$, i.e., there is an embedding $\phi_{\zeta}:[0,1]\rightarrow\mathcal{Q}$ such that $\phi_{\zeta}([0,1])=\pi(\zeta(\mathbb{R}))$. Instead, $\zeta(t)$ is a weak geometric eigenmode if $\phi_{\zeta}:[0,1]\rightarrow\mathcal{Q}$ is an immersion and $\phi_{\zeta}(0)\neq\phi_{\zeta}(1)$. The points $\phi_{\zeta}(0)$ and $\phi_{\zeta}(1)$ are called brake points [63].
Definition 9 ((Weak) Eigenmanifold).
An LSM $\mathcal{M}\subseteq T^{*}Q$ is an (weak) Eigenmanifold if all trajectories contained within it are (weak) geometric eigenmodes.
Definition 10 (Generator).
We call the set $\mathcal{R}\subset\mathcal{M}$ of brake points the generator of $\mathcal{M}$.
The generator satisfies that $\mathcal{M}=\Psi^{\mathbb{R}}_{X_{H}}(\mathcal{R})$, i.e., it generates $\mathcal{M}$ by forward evolution of the dynamics $X_{H}$.
Prior definitions of Eigenmanifolds ([22, Sec. 7]) explicitly assumed the existence of the generator. We conclude from Theorem 7 that this assumption is not necessary, the generator exists for any family of weak geometric eigenmodes.
A lower bound for the number of Eigenmanifolds follows from Corollary 2 and Theorem 7:
Corollary 11 (Number of Eigenmanifolds).
Consider a system of the type (5), (6), with an equilbrium $\bar{q}$ a minimum of the potential $V(q)$. If $m\leq n$ of the $n$ eigenvalues $\omega^{2}_{k}$ of the $(1,1)$-tensor $M(\bar{q})^{-1}\frac{\partial^{2}V}{\partial q^{T}\partial q}(\bar{q})$ are mutually non-resonant, and non-resonant with the remaining eigenvalues, then there exist, locally, at least $m$ unique Eigenmanifolds.
4.2 Time and spatially symmetric Lyapunov subcenter manifolds
Here, we show subjecting the Lyapunov subcenter manifold to an additional spatial symmetry causes all trajectories to pass through the equilibrium configuration.
With $\bar{x}=\arg\min V(x)\in\mathcal{Q}$ the equilibrium configuration of (2), let $\varphi:\mathcal{Q}\rightarrow\mathcal{Q}$ be a diffeomorphism such that:
$\varphi(\bar{x})$ $=\bar{x}\,$, (34)
$\varphi\circ\varphi$ $=\text{id}_{\mathcal{Q}}\,$, (35)
$\varphi_{*}(\bar{x})$ $=-\text{id}_{T_{\bar{x}}\mathcal{Q}}\,.$ (36)
We denote a class of equivariant symmetries by $\big(x,P,t\big)\rightarrow\big(\varphi(x),(\varphi^{-1})^{*}P,t\big)$, which we write as $(\sigma_{2},\tau_{2})$, with
$\sigma_{2}((x,P))=(\varphi(x),(\varphi^{-1})^{*}P)\,,\quad\tau_{2}=1\,.$ (37)
Symmetries of this class apply to the CM (2) with Hamiltonian (1) when $M=\varphi^{*}M$ and $V=V\circ\varphi$, as we show in Appendix A.2.
A local chart exists (see Appendix A.2) that maps $\bar{x}$ to $q=0$, and where the symmetry reads $(q,p,t)\rightarrow(-q,-p,t)$, such that $(\sigma_{2},\tau_{2})$ can be written as $(S_{2},\tau_{2})$ with
$S_{2}((q,p))=(-q,-p)\,,\quad\tau_{2}=1\,.$ (38)
We refer to these coordinates as $\varphi$-equivariant coordinates, and in these coordinates the conditions $M=\varphi^{*}M$ and $V=V\circ\varphi$ read
$M(q)$ $=M(-q)\,$, (39)
$V(q)$ $=V(-q)\,.$ (40)
Theorem 12 (Rosenberg manifolds).
Given the conditions of Theorem 4, and additionally $\varphi:\mathcal{Q}\rightarrow\mathcal{Q}$ satisfying (34) to (36) and
  1. (i)
    $V=V\circ\varphi$ and $M=\varphi^{*}M$
  2. (ii)
    The equilibrium $\bar{x}$ is the only fixed point of $\varphi$ such that $V(\bar{x})\leq V(Q)$ for $Q\in\mathcal{M}$
then a unique LSM $\mathcal{M}\subset T^{*}\mathcal{Q}$ associated with the eigenspace $E_{0}\subset T_{(\bar{q},0)}T^{*}\mathcal{Q}$ has, in addition to the properties in Theorem 7, the properties
  1. 1.
    Equilibrium: all trajectories $\zeta(t)\in\mathcal{M}$ pass through the equilibrium configuration $\bar{x}\in\pi(\zeta(\mathbb{R}))$.
  2. 2.
    Time to equilibrium: with $T_{\zeta}$ the period of $\zeta(t)$, then the points with $\pi(\zeta(t))=\bar{x}$ are a time $T_{\zeta}/2$ apart and a time $T_{\zeta}/4$ from the brake points.
{pf}
See Appendix A.3. We make two definitions based on Theorem 12:
Definition 13 ((Weak) Geometric Rosenberg Mode).
A $T$-periodic oscillation $\zeta:\mathbb{R}\rightarrow T^{*}\mathcal{Q}$ is a (weak) geometric Rosenberg mode with respect to an equilibrium $\bar{x}$ if it is both a (weak) geometric eigenmode, and $\bar{x}\in\pi_{Q}(\zeta(\mathbb{R}))$, i.e., if $\zeta$ passes through the equilibrium configuration $\bar{x}$ for some momentum $P$.
Definition 14 ((Weak) Rosenberg Manifold).
An LSM $\mathcal{M}\subseteq T^{*}Q$ is a (weak) Rosenberg manifold if all trajectories contained within it are (weak) Rosenberg modes.
Remark 15.
Definitions 13 and 14 of geometric Rosenberg modes and Rosenberg manifolds align with the definitions put forth in [22], but disagree with other definitions in the literature: originally [18] defined Rosenberg modes as modes for which the coordinates $q_{1}(t),\cdots,q_{n}(t)$ oscillate in unison, necessarily passing through the unique equilibrium at $q=0$. This is ill-defined in a coordinate-free setting: such coordinates can be found for any geometric eigenmode, and the property is not invariant under changes of coordinates. In Definition 13, instead, there is no reference to oscillation in unison. More recently [19] defined extended Rosenberg modes, which allow for coordinates of periodic motions to oscillate asynchronously. This definition can be read to include arbitrary periodic modes, geometric eigenmodes and extended Rosenberg modes, making it non-specific. Definition 13 excludes arbitrary periodic oscillations and makes geometric Rosenberg modes a subset of geometric eigenmodes.
$\mathcal{Q}$
$\bar{x}$
$\left(x,0\right)$
$\left(x(t),P(t)\right)$
(a) Weak geometric eigenmode
$\mathcal{Q}$
$\bar{x}$
$\left(x,0\right)$
$\left(\varphi(x),0\right)$
$\left(x(t),P(t)\right)$
(b) Weak geometric Rosenberg mode
Figure 2: Example of a weak geometric eigenmode in panel (a) and a $\varphi$-equivariant weak geometric Rosenberg mode in panel (b). Extending their non-weak counterparts, these allow for self-intersections in configuration space.
5 Examples
This section presents two examples. The first example highlights various aspects of Theorems 7 and 12 for systems with $\text{dim}(\mathcal{Q})=2$, to ease visual understanding. The second example is for $\text{dim}(\mathcal{Q})=5$, and investigates the conditions M(q)=M(q)M(q)=M(-q), V(q)=V(q)V(q)=V(-q) of Theorem 12 for a non-trivial system.
5.1 Various 2-DoF systems
We consider the system of coupled masses in Fig. 3(a), which is a Euclidean system with $\mathcal{Q}_{C}=\mathbb{R}^{2}$ and a curvature-free metric tensor $M_{C}$ given by
$M_{C}(q)=\begin{bmatrix}m&0\\ 0&m\end{bmatrix}\,$, (41)
and the double pendulum with parallel elasticity shown in Fig. 3(b), which is a non-Euclidean system with $\mathcal{Q}_{DP}=T^{2}$ and metric tensor $M_{DP}$ given by
$M_{DP}(q)=\begin{bmatrix}I+3d^{2}m(1+2\cos(q_{2})&d^{2}m(1+\cos(q_{2}))\\ d^{2}m(1+\cos(q_{2}))&I+d^{2}m\end{bmatrix}\,$. (42)
$m$
$m$
$q_{1}$
$q_{2}$
(a) Euclidean system
$q_{1}$
$q_{2}$
(b) Non-Euclidean system
Figure 3: Examples with $\dim(\mathcal{Q})=2$: the system of two masses in panel (a) has a Euclidean configuration space, $\mathbb{R}^{2}$ with a constant inertia tensor. The double pendulum in panel (b) has a non-Euclidean configuration space, $T^{2}$ with a non-zero curvature, where the associated inertia tensor is non-constant in any choice of coordinates.
We examine the effect of different potential functions, see also Fig. 4:
$V_{s1}(q)$ $=\nicefrac{{1}}{{2}}k(q_{2})^{2}-dmg(2\cos(q_{1})+\cos(q_{1}+q_{2})),$
$V_{s2}(q)$ $=\nicefrac{{1}}{{2}}kq_{1}^{2}+\nicefrac{{1}}{{2}}k(q_{2}-\nicefrac{{\pi}}{{2}})^{2}\,$,
$V_{a}(q)$ $=\nicefrac{{1}}{{2}}k(q_{2}-\nicefrac{{\pi}}{{2}})^{2}-dmg(2\cos(q_{1})+\cos(q_{1}+q_{2}))\,.$
$k$
$g$
(a) $V_{s1}$
$k$
$k$
(b) $V_{s2}$
$g$
$k$
(c) $V_{a}$
Figure 4: Equilibrium configurations of the three tested potentials: $V_{s1}$ with nonlinear terms from gravity and $V_{s2}$ with only quadratic spring terms are symmetric under $(S_{2},\tau_{2})$, in the appropriate coordinates. $V_{a}$ is obtained from $V_{s1}$ by shifting the equilibrium of the spring to $\nicefrac{{\pi}}{{2}}$ such that $V_{a}(\bar{q}+q)\neq V_{a}(\bar{q}-q)$ for the minimum $\bar{q}$ of $V_{a}$, making $V_{a}$ not symmetric.
Parameters are chosen as $m=0.4\text{kg},I=\nicefrac{{1}}{{12}}\text{kgm}^{2},d=1\text{m},k=10\frac{\text{Nm}}{\text{rad}},g=9.81\text{m}/\text{s}^{2}$.
By different pairings of the above metric tensors and potential functions, we obtain a variety of systems with dynamics governed by (5), (6)., subsequently denoted by a tuple $(M,V)$. The system $(M_{C},V_{s2})$ is linear, and we exclude it from the results. The results from numerical continuation are summarized in Figs. 5, 6 and in Tab. 2. All systems satisfy the conditions of Theorem 7 for the existence of two unique Eigenmanifolds, and show the existence of two Eigenmanifolds, as expected.
The double pendulum $(M_{DP},V_{s1})$ and the system $(M_{C},V_{s1})$ additionally satisfy the conditions of Theorem 12, and Rosenberg manifolds are found. For $(M_{DP},V_{s1})$, example Rosenberg modes and the generators are shown in Figs. 5(a) & 5(b). For the system $(M_{C},V_{s1})$, this is shown in Fig. 5(h) & 5(i).
Table 2: Summary of tested 22\--DoF examples
System System type Potential Inertia Thm. 12 satisfied? Figure
$(M_{DP},V_{s1})$ non-Euclidean Symmetric Symmetric Fig. 5(a), 5(b)
$(M_{DP},V_{a})$ non-Euclidean Asymmetric Symmetric Fig. 5(d), 5(e)
$(M_{DP},V_{s2})$ non-Euclidean Symmetric Asymmetric Fig. 5(f), 5(g)
$(M_{C},V_{s1})$ Euclidean Symmetric Symmetric Fig. 5(h), 5(i)
$(M_{C},V_{a})$ Euclidean Asymmetric Symmetric Fig. 5(j), 5(k)
$q_{2}$ in rad
$q_{1}$ in rad
(a) $(M_{DP},V_{s1}),R_{1}$
$q_{2}$ in rad
$q_{1}$ in rad
(b) $(M_{DP},V_{s1}),R_{2}$
Generator $R_{11}$
Generator $R_{21}$
Generator $R_{12}$
Generator $R_{22}$
Example Trajectory
(c) Legend
$q_{2}$ in rad
$q_{1}$ in rad
(d) $(M_{DP},V_{a}),R_{1}$
$q_{2}$ in rad
$q_{1}$ in rad
(e) $(M_{DP},V_{a}),R_{2}$
$q_{1}$ in rad
$q_{2}$ in rad
(f) $(M_{DP},V_{s2}),R_{1}$
$q_{2}$ in rad
$q_{1}$ in rad
(g) $(M_{DP},V_{s2}),R_{2}$
$q_{2}$ in rad
$q_{1}$ in rad
(h) $(M_{C},V_{s1}),R_{1}$
$q_{2}$ in rad
$q_{1}$ in rad
(i) $(M_{C},V_{s1}),R_{2}$
$q_{2}$ in rad
$q_{1}$ in rad
(j) $(M_{C},V_{a}),R_{1}$
$q_{2}$ in rad
$q_{1}$ in rad
(k) $(M_{C},V_{a}),R_{2}$
Figure 5: Results of numerical continuation for different example systems generated by combining different potential functions and inertia tensors. The solid lines show computed generators of the system. The pair of blue and orange solid lines show the two generators associated to the first Eigenmanifold and the pair of purple and red the two associated to the second Eigenmanifold. The dashed blue line shows an example modal oscillation of the system for one energy. The conditions stated by Theorem 12 are satisfied for the cases (a,b) and (h,i).
The systems that do not satisfy Theorem 12 are the double pendulum with asymmetric potential $(M_{DP},V_{a})$ in Figs. 5(d) & 5(e), the double pendulum with symmetric potential $(M_{DP},V_{s2})$ in Figs. 5(f) & 5(g), and the two-mass system with asymmetric potential $(M_{C},V_{a})$ in Figs. 5(j) & 5(k) do not satisfy Theorem, such that most modes are no continuous families of Rosenberg modes. However, we observe that isolated geometric Rosenberg modes on Eigenmanifolds can be found, which are not excluded by Theorem 12. The isolated modes are more easily recognized in Fig. 6, which shows the minimum potential energy $V$ along a given mode together with the energy of the starting point of that mode for all five systems. Whenever this minimum potential is zero, the corresponding mode necessarily passes through the equilibrium. Thus, isolated zeros of this function identify isolated geometric Rosenberg modes within Eigenmanifolds.
Finally, although Theorem 12 provides only a sufficient condition for Rosenberg manifolds, we could not find any example of Rosenberg’s manifolds for systems not fulfilling the hypotheses of Theorem 12.
Minimum potential along mode in J
Energy of starting point in J
Generator $R_{11}$
Generator $R_{21}$
Generator $R_{12}$
$(M_{DP},V_{s_{1}})$
$(M_{DP},V_{a})$
$(M_{C},V_{s_{1}})$
$(M_{DP},V_{s_{2}})$
$(M_{C},V_{a})$
(a) Generators $R_{1}$
Minimum potential along mode in J
Energy of starting point in J
Generator $R_{11}$
Generator $R_{21}$
Generator $R_{12}$
$(M_{DP},V_{s_{1}})$
$(M_{DP},V_{a})$
$(M_{C},V_{s_{1}})$
$(M_{DP},V_{s_{2}})$
$(M_{C},V_{a})$
(b) Generators $R_{2}$
Figure 6: Minimal potential energy for modes of the example systems, split by modes collected on first and second generator.
5.2 Example of a high dimensional system
We analyze a more complex system satisfying the symmetry conditions required by Theorem 12, and find families of geometric Rosenberg modes. We take a quintuple pendulum, see Figure 8(a), set all masses to $m=0.4\mathrm{kg}$ and set the link lengths to $l=1.0\mathrm{m}$. For the potential, we use a diagonal stiffness matrix $K=20I_{5}\frac{\mathrm{Nm}}{\mathrm{rad}}$ for a spring term and set $g=-9.81\frac{\mathrm{m}}{\mathrm{s}^{2}}$. As coordinates $q$, we choose joint-angles in radians, with the equilibrium configuration $q=0$ corresponding to straight-down reference configuration. We do not report the full inertia tensor $M(q)$, but remark that M(q)=M(q)M(q)=M(-q) and V(q)=V(q)V(q)=V(-q) for the chosen $q$.
The eigenvalues of $M(0)^{-1}\frac{\partial^{2}V}{\partial q\partial q}(0)$ are given by $(1123.58,787.417,255.266,103.333,56.6527)$, which are mutually non-resonant. Thus, Corollary 11 predicts the existence of 5 weak Eigenmanifolds, and by Theorem 12 these are geometric Rosenberg manifolds. To numerically show this, we compute the Eigenmanifolds of the quintuple pendulum up to an energy level of $E_{max}=100\mathrm{J}$. The lower triangular matrix in Fig. 7(a) shows the projections of the corresponding generators onto $q_{i}q_{j}$-planes. For each generator, we highlight the configuration of maximal potential energy, which is indicated by the dots in the figure. When we take these configurations as initial configuration for simulating the pendulum. The resulting oscillations pass through the equilibrium configuration and classify as geometric Rosenberg modes. Projections of these trajectories are shown in the upper triangular matrix in Fig. 7(a).
Generators
$q_{1}$
$q_{2}$
$q_{3}$
$q_{4}$
$q_{5}$
$q_{1}$
$q_{2}$
$q_{3}$
$q_{4}$
$q_{5}$
$q_{1}$
$q_{2}$
$q_{3}$
$q_{4}$
$q_{5}$
$q_{1}$
$q_{2}$
$q_{3}$
$q_{4}$
$q_{5}$
High Energy
Modal
Oscillation
$R_{11}$
$R_{12}$
$R_{21}$
$R_{22}$
$R_{31}$
$R_{32}$
$R_{41}$
$R_{42}$
$R_{51}$
$R_{52}$
(a) Generators and high energy modal oscillations of the quintuple pendulum shown in configuration space. The lower triangular matrix of plots shows the generators projected onto different $q_{i}q_{j}$-planes. The upper triangular matrix of plots shows the corresponding modal oscillation for the highest energy on the color-matching generator. The system satisfies Theorem 12, so all modal oscillations pass through the equilibrium.
Generators
One of the Modal Oscillations
Generator 1
Generator 2
Generator 3
Generator 4
Generator 5
$x$
$y$
$y$
(b) Cartesian version of Fig. 7(a). The top row shows the generators as Cartesian paths of the joints and the bottom row shows one modal oscillation. All modal oscillation pass through the equilibrium, in which all links of the pendulum point straight down.
Figure 7: Generators of a quintuple pendulum.
We only show the generators and highest energy modal oscillation in Fig. 7(b), but report that the modal oscillation for all energy levels pass through the equilibrium configuration. We also show various modal oscillations on the fourth generator, in Fig. 8(b). Again, we observe that all modal oscillation pass through a configuration where all links point straight downwards.
$q_{3}$
$q_{4}$
$q_{5}$
$k_{1}$
$k_{2}$
$k_{3}$
$k_{4}$
$k_{5}$
$g$
$l$
$l$
$l$
$l$
$l$
$m$
$m$
$m$
$m$
$m$
$q_{1}$
$q_{2}$
(a) Quintuple elastic pendulum with gravity
$E=6.05J$
$E=26.16J$
$E=52.41$
$E=71.77$
$E=75.46$
$x$
$x$
$y$
(b) Modes of different energy levels on the fourth Rosenberg manifold, in Cartesian space. The left shows a number of initial configurations, and the right highlights the modal oscillations of the tip.
Figure 8: Quintuple pendulum and various Rosenberg modes.
6 Conclusion
Lyapunov subcenter manifold (LSM) theory shows that existence and uniqueness of families of periodic orbits follows from non - resonance conditions. For conservative mechanical systems, we showed that non - resonance conditions on the inertia tensor and potential energy guarantee that LSMs have stronger properties properties of weak Eigenmanifolds, in which all orbits are periodic brake trajectories. When the configuration space is two - dimensional, unique LSMs are Eigenmanifolds, in which the configuration trajectory of modes does not self-intersect. We proved that the generator, the collection of brake points on an Eigenmanifold, is a uniquely defined 1D connected submanifold. We also presented conditions on the Riemannian inertia tensor and potential energy that results in unique Eigenmanifolds to fulfill the stronger properties of Rosenberg manifolds. In numerical examples, the absence of such continuous families was observed when these conditions were violated. The results confirm the validity of the presented Theorems, and present a first step towards stronger properties of nonlinear normal modes as practically feasible target trajectories.
Appendix A Proofs
A.1 Time-symmetric Lyapunov subcenter manifolds
We prove Theorem 7 in steps, treating the properties 1 through 4 in order.
A.1.1 Preliminary: brake points in time-symmetric Hamiltonian systems
We begin by repeating a well-known theorem on orbits in time-symmetric Hamiltonian systems.
Theorem 16 (Theorem 4.1, [61]).
Given Hamiltonian dynamics with H(x,P)=H(x,P)H(x,P)=H(x,-P).
  1. (a)
    An evolution $\zeta(t)=(x(t),P(t))$ of (2) satisfies
    $\zeta(\mathbb{R})={\sigma}_{1}(\zeta(\mathbb{R}))\,.$ (43)
    if and only if there is $t^{\prime}$ such that $P(t^{\prime})=0$.
  2. (b)
    Given $T>0$, a $T$-periodic orbit $(x(t),P(t))=(x(t+T),P(t+T))$ encounters either 0 or 2 distinct points with $P(t)=0$.
{pf}
This directly follows from Theorem 4.1a in [61] with $\text{Fix}(\sigma_{1})=\mathcal{Q}\times 0\subset T^{*}\mathcal{Q}$, i.e., the fixed set of the symmetry $\sigma_{1}$ is the zero section of $T^{*}\mathcal{Q}$. By Theorem 16a, any trajectory $(x(t),P(t))$ with $P(0)=0$ satisfies
x(t)=x(t).x(t)=x(-t)\,. (44)
i.e., the backwards and forwards evolution of the configuration from a starting point $(x(0),0)$ are identical. Braking points appear to reflect $x(t)$. If the orbit is also periodic, $x(t)$ that satisfies (44) must reflects at a second braking point. Theorem 16b shows that there can be no further reflections in between.
A.1.2 Period between brake points
We show that the period and the time between brake points are related: given a period $T_{\zeta}$, the time to go from one brake point to another is $\Delta t=T_{\zeta}/2$.
Lemma 17.
Given an evolution $\zeta(t)=(x(t),P(t))$ of (2). If there are two distinct points $(x_{1},0)$, $(x_{2},0)\in\zeta(\mathbb{R})$ and denoting $\Delta t=\min\{t>0\,|\,(x_{2},0)=\Psi^{t}_{X_{H}}(x_{1},0)\}$, then the evolution is periodic with period $T_{\zeta}=2\Delta t$.
{pf}
Let $\zeta(t)$ be such that $\zeta(0)=(x_{1},0)$ and $\zeta(\Delta t)=(x_{2},0)$. We show that $\zeta(-\Delta t)=\zeta(\Delta t)$, such that the evolution must be periodic with $T_{\zeta}=2\Delta t$. Let
$\bar{\zeta}(t)={\sigma}_{1}\zeta(\tau_{1}t)=\big(x(-t),-P(-t)\big)$ (45)
Substituting directly into (45):
$\bar{\zeta}(-\Delta t)=(x(\Delta_{t}),-P(\Delta_{t}))=(x_{2},0)\,.$ (46)
However, $\bar{\zeta}(0)=(x_{1},0)$ and it follows from uniqueness of solutions that $\zeta(t)=\bar{\zeta}(t)$, so $\zeta(-\Delta t)=\bar{\zeta}(-\Delta t)=\zeta(\Delta t)$, as required.
A.1.3 Immersion and brake point properties
Lemma 18.
Given $T>0$, a $T$-periodic orbit $\zeta(t)$ there is an immersion $\phi_{\zeta}:[0,1]\rightarrow\mathcal{Q}$ satisfying (33) if and only if $\zeta(t)$ satisfies the symmetry
$\zeta(\mathbb{R})={\sigma}_{1}(\zeta(\mathbb{R}))\,.$ (47)
{pf}
If: given the symmetry, Theorem 16a guarantees a point with $P(t^{\prime})=0$. Then periodicity and Theorem 16b, Lemma 17 guarantee that $P(t^{\prime}+T/2)=0$. Then the segment $x([t^{\prime},t^{\prime}+T/2])\subset\mathcal{Q}$ is such that
$x([t^{\prime},t^{\prime}+T/2])=x([t^{\prime}+T/2,t^{\prime}+T])\,.$ (48)
The map
$\phi_{\zeta}(s)=\pi_{\mathcal{Q}}(\zeta(t^{\prime}+sT/2))$ (49)
is an immersion mapping the interval $[0,1]$ to $x([t^{\prime},t^{\prime}+T/2])$, where $\frac{\text{d}}{\text{d}s}\phi_{\zeta}(s)\neq 0$ for $s\in(0,1)$ since $P\neq 0$, and at $s=0,s=1$, we have that $\phi_{\zeta}(0)=x(t^{\prime})$ and $\phi_{\zeta}(1)=x(t^{\prime}+T/2)$ as required. Since, $x([t^{\prime},t^{\prime}+T/2])=x([t^{\prime}+T/2,t^{\prime}+T])$, this extends to
$\phi_{\zeta}(s)=\pi(\zeta(\mathbb{R}))\,$, (50)
and $\zeta(t)$ is a weak geometric eigenmode.
Only if: given a periodic orbit that for which $\phi_{\zeta}(s)$ can be found, there are necessarily two points with $P(t)=0$, and the symmetry is satisfied as a consequence of Theorem 16.
Theorem 19 (Time-symmetric modes).
Given the conditions of Theorem 4, all $\zeta(t)$ in the unique LSM $\mathcal{M}\subset T^{*}\mathcal{Q}$ are such that
$\zeta(\mathbb{R})={\sigma}_{1}(\zeta(\mathbb{R}))\,.$ (51)
{pf}
The symmetry $(S_{1},\tau_{1})$ satisfies the conditions of Theorem 5:
$S_{1}((0,0))=(0,0)\,$, (52)
and the eigenspace $E_{0}:=D\oplus M(\bar{x})D$ is such that
$E_{0}=\frac{\partial S_{1}}{\partial z}E_{0}=\begin{bmatrix}I&0\\ 0&-I\end{bmatrix}E_{0}\,.$ (53)
Thus, the LSM $\mathcal{M}$ is such that
$\mathcal{M}$ $=\sigma_{1}(\mathcal{M})\,$, (54)
$\zeta(\mathbb{R})$ $=\sigma_{1}\big(\zeta(\mathbb{R})\big)\,.$ (55)
Then the immersion property of Theorem 7 is a direct consequence of Theorem 19 and Lemma 18. The brake point property of Theorem 7 is a direct consequence of Theorem 19 and Lemma 17.
A.1.4 Embedding Property
Theorem 20 (Embedding Property).
Given the conditions of Theorem 4 and the unique LSM $\mathcal{M}\subset T^{*}\mathcal{Q}$, there is a neighborhood $\mathcal{M}^{\prime}\subset\mathcal{M}$ containing the equilibrium where there is an embedding $\phi_{\zeta}:[0,1]\rightarrow\mathcal{Q}$ satisfying (33). When $\dim\mathcal{Q}=2$, then $\mathcal{M}^{\prime}=\mathcal{M}$.
{pf}
By the immersion property 7, an immersion $\phi_{\zeta}$ is already guaranteed. To be an embedding, $\phi_{\zeta}$ must further be injective, i.e., $x(t)$ should not self-intersect. The limiting behavior of $\zeta(t)=\Psi^{t}_{X_{H}}(\zeta_{0})\in\mathcal{M}_{0}$ as $\zeta_{0}=(x,0)$ approaches $(\bar{x},0)$ is that of a linear oscillation, which necessarily does not self-intersect. Hence, there must also be a neighborhood $\mathcal{M}^{\prime}_{0}\subset\mathcal{M}$ where $\phi_{\zeta}$ is an embedding.
We now cover the case $\dim\mathcal{Q}=2$. One of two scenarios must occur in transitioning from $\mathcal{M}^{\prime}$ to $\mathcal{M}$ (see also Figures 9(a) and 9(b)).
$\mathcal{Q}$
$\bar{x}$
$\left(x_{2},0\right)$
$\left(x_{1},0\right)$
$\left(x(t),P(t)\right)$
(a) First order self-intersection of a geometric eigenmode.
$\mathcal{Q}$
$\bar{x}$
$\left(x,P\right)$
$\left(x(t),P(t)\right)$
(b) Higher order self-intersection of a geometric eigenmode.
Figure 9: Possible self-intersections of Eigenmodes along a generator, assuming a two dimensional configuration space. In transitioning from a geometric eigenmode to a weak geometric eigenmode, either (a) a first order intersection takes places or (b) a higher order intersection takes place.
Scenario a, a first order intersection where either one of the end-points $x_{1}$ or $x_{2}$ passes through a $x_{3}$ contained within $\pi_{\mathcal{Q}}\big(\zeta(\mathbb{R})\big)$. A first order intersection would imply e.g., $V(x_{3})=V(x_{2})$ and by conservation of energy the momentum at $x_{3}$ must be $P_{3}=0$. By Theorem 16b, such a third point with $P=0$ on the same mode $\zeta(t)$ is not possible.
Scenario b concerns a higher order intersection where $\zeta(\mathbb{R})$ is tangent to itself at some point $(x,P)\in\zeta(\mathbb{R})$ that is not an end-point. Uniqueness of solutions forbids this transition, since the system must have a unique solution through any $(x,P)$.
Since the transition from embedded to immersed-but-not-embedded modes is impossible for $\dim\mathcal{Q}=2$, it must be that all modes are embedded and $\mathcal{M}^{\prime}=\mathcal{M}$ for this case.
Remark 21.
The proof breaks for $\dim\mathcal{Q}=2$ down in higher dimensions, since other first order intersections of points $(x,P)\in\zeta(\mathbb{R})$ that are end-points become possible. However, we expect that self-intersections become less “likely” with increasing $n$, being most common for $\text{dim}(\mathcal{Q})=3$.
A.1.5 Generator Property
To show that the generator property holds, we prove the following theorem:
Theorem 22 (Existence of generator).
Given a unique LSM $\mathcal{M}\subset T^{*}\mathcal{Q}$ that satisfies the immersion property in Theorem 7. Then the set $\mathcal{R}:=\{(x,0)\,|\,(x,0)\in\mathcal{M}\}$ is a connected, 1D submanifold containing the equilibrium.
{pf}
We first show that $\mathcal{R}=\{(x,0)\,|\,(x,0)\in\mathcal{M}\}$ is a regular 1-dimensional submanifold of $\mathcal{M}$. We use the regular level set theorem [64, Corr. 5.24], which states that $\mathcal{R}\subset\mathcal{M}$ is a closed, embedded submanifold of dimension 1 if there is a function $\alpha:\mathcal{M}\rightarrow\mathbb{R}$ such that $\mathcal{R}=\alpha^{-1}(0)$ and $\text{Im}(\text{d}\alpha)=T_{0}\mathbb{R}$. To this end, note that
$\alpha(x,P)=M^{-1}(\text{d}V(x),P)+M^{-1}(P,P)\,.$ (56)
is as required: points on the generator satisfy that $P=0$, which corresponds to $\alpha(x,0)=0$. For $\text{Im}(\text{d}\alpha)=T_{0}\mathbb{R}$, note that $\mathcal{M}$ is tangent to the flow $X_{H}$, and apply $\text{d}\alpha(x,P)$ to the vector $X_{H}(x,P)$ at $(x,0)\in\mathcal{M}$:
$\text{d}\alpha(X_{H})$ $=M^{-1}(\text{d}V(x),\dot{P})$ (57)
$=M^{-1}(\text{d}V(x),\text{d}V(x))\neq 0\,.$
Thus, it follows from the regular level set theorem [64] that $\mathcal{R}$ is a closed, embedded 1D submanifold. This holds where $\text{d}V(x)\neq 0$, i.e., at all points other than the equilibrium. The equilibrium itself may be included as a limit point, since geometric eigenmodes arbitrarily close to $(\bar{x},0)$ may be found by definition of $\mathcal{M}$.
From the closedness of $\mathcal{R}$ it can be shown that it is connected: Consider coordinates $X:\mathcal{M}\rightarrow\mathbb{R}^{2}$, such that orbits are mapped to lines of constant radius, and angle $\theta$ ranges from 0 to $2\pi$ over a full oscillation. Then the generator is parameterized as
$\mathcal{R}=\big(r(z),\theta(z)\big)\,.$ (58)
Let the coordinates be such that the antipodal brake points guaranteed by Lemma 17 are
$\sigma_{1}(\mathcal{R})=\big(r(z),-\theta(z)\big)\,.$ (59)
Assume, for the sake of contradiction, that $\mathcal{R}$ is not connected but consists of connected, closed submanifolds $\mathcal{R}_{1},\mathcal{R}_{2}\subset\mathcal{R}$. Let $\bar{r}_{1}$ be the radius on the boundary of $\mathcal{R}_{1}$. By closure, $\bar{r}_{1}$ must also lie on the boundary of $\mathcal{R}_{2}$. Yet, if $\mathcal{R}_{0,i}$, $\mathcal{R}_{0,j}$, $\sigma_{1}(\mathcal{R}_{0,i})$ and $\sigma_{1}(\mathcal{R}_{0,j})$ are distinct for $\bar{r}_{1}$, there would have to be geometric eigenmodes with four brake points. This contradicts Theorem 16b, so the generator must be connected.
A.2 Spatial symmetry
We show that a set of equivariant coordinates exists in which $\bar{q}=0$ and $\varphi:\mathcal{Q}\rightarrow\mathcal{Q}$ is the map $q\mapsto-q$ and $S_{2}(q,p)=(-q,-p)$. Afterwards, we use this set of coordinates to prove the symmetry of the equations.
A.2.1 Equivariant coordinates
Given an diffeomorphism $\varphi:\mathcal{Q}\rightarrow\mathcal{Q}$, that leaves an equilibrium $\bar{x}\in\mathcal{Q}$ fixed, satisfies that $\varphi_{*}(\bar{q})=-\text{id}_{T_{\bar{q}}\mathcal{Q}}$ and $\varphi\circ\varphi=\text{id}_{\mathcal{Q}}$.
Pick any chart $(U,X)$ with chart-region $\bar{x}\in U\subset\mathcal{Q}$ and (diffeomorphic) chart-map $X:\mathcal{Q}\rightarrow\mathbb{R}^{n}$. We want to find a chart $(V^{\prime},Y)$ such that $S_{2}(q,p)=(-q,-p)$. For this it is sufficient that $Y\circ\varphi\circ Y^{-1}(q)=-q$, or equivalently
$Y\circ\varphi=-Y\,.$ (60)
Define a candidate chart $(V,Y)$ by
$V$ $:=\{x\in U\,|\,\varphi(x)\in U\}\,$, (61)
$Y$ $:=\frac{1}{2}(X-X\circ\varphi)\,.$ (62)
Then the coordinate description of the symmetry is $Y\circ\varphi(x)=-Y(x)$, as desired. Also note that $V$ is well defined since $\varphi\circ\varphi=\text{id}_{\mathcal{Q}}$.
It remains to be shown that $Y$ is a valid chart-map, i.e., that $Y$ is a local diffeomorphism. This follows from the inverse function theorem [65, Theorem 1.3.12] by showing that $Y_{*}(\bar{x})$ is full rank:
$Y_{*}(\bar{x})=\frac{1}{2}(X_{*}(\bar{x})-X_{*}(\bar{x})\varphi_{*}(\bar{x}))=X_{*}(\bar{x})\,.$ (63)
Here, the first step uses that $\varphi(\bar{x})=\bar{x}$ and the second step uses that $\varphi_{*}(\bar{x})=-\text{id}_{T_{\bar{x}}\mathcal{Q}}$. This is indeed full rank, since $X$ is itself a local diffeomorphism. Thus, a neighborhood $V^{\prime}\subset V$ of $\bar{x}$ exists such that $(V^{\prime},Y)$ is a valid chart with the desired properties.
A.2.2 Proof of spatial symmetry
The spatial symmetry is stated in $\varphi$-equivariant coordinates (see also Appendix A.2.1).
Theorem 23 (Spatial symmetry).
If $M=\varphi^{*}M$ and $V=V\circ\varphi$ hold for $\varphi$ satisfying (34) to (36), then the symmetry
$\sigma_{2}((x,P))=(\varphi(x),(\varphi^{-1})^{*}P)\,,\quad\tau_{2}=1\,.$ (64)
satisfies
$X_{H}\circ\sigma=\tau\sigma_{*}X_{H}\,.$ (65)
{pf}
We use $\varphi$-equivariant coordinates. Then $M=\varphi^{*}M$ and $V=V\circ\varphi$ read, respectively:
$M(q)$ $=M(-q)\,$, (66)
$V(q)$ $=V(-q)\,$, (67)
I.e., $V(q)$ and the component-functions $M(q)$ are even functions. The dynamics (2) read
$f\big(z\big)=\begin{pmatrix}M(q)^{-1}p\\ -\frac{\partial}{\partial q}(p^{\top}M(q)^{-1}p)-\frac{\partial V}{\partial q}(q)\end{pmatrix}\,$. (68)
And condition 65 reads
$\frac{\partial S_{2}}{\partial z}f\big(z\big)=f\big(S_{2}(z)\big)\,.$ (69)
Note that the gradient of an even function h(q)=h(q)h(q)=h(-q) is odd, i.e., $\frac{\partial h}{\partial q}(q)=-\frac{\partial h}{\partial q}(-q)$. Thus the right-hand side is
$f\big(S_{2}(z)\big)$ $=f(-q,-p)$
$=\begin{pmatrix}-M(-q)^{-1}p\\ -1/2\frac{\partial}{\partial s}(p^{\top}M(s)^{-1}p)_{s=-q}-\frac{\partial V}{\partial s}_{s=-q}\,.\end{pmatrix}$
$=\begin{pmatrix}-M(q)^{-1}p\\ \frac{\partial}{\partial q}(p^{\top}M(q)^{-1}p)+\frac{\partial V}{\partial q}\,.\end{pmatrix}\,$. (70)
For the left-hand side we get
$\tau_{1}\frac{\partial S_{2}}{\partial z}F\big(z\big)$ $=\begin{bmatrix}-I&0\\ 0&-I\end{bmatrix}F(z)$
$=\begin{pmatrix}-M(q)^{-1}p\\ \frac{\partial}{\partial q}(p^{\top}M(q)^{-1}p)+\frac{\partial V}{\partial q}(q)\,.\end{pmatrix}\,$. (71)
Both sides are the same, so the symmetry condition 65 holds. This completes the proof.
In the same $\varphi$-equivariant coordinates, the symmetry means that if a trajectory $\big(q(t),p(t)\big)$ is a solution of the dynamics (5), (6) then also $\big(-q(t),-p(t)\big)$ is a solution.
A.3 Time- and spatially symmetric Lyapunov subcenter manifolds
In this section, the properties guaranteed by Theorem 12 are proven in order.
Theorem 24 (Spatially symmetric modes).
Given the conditions of Theorem 4, and if $\varphi:\mathcal{Q}\rightarrow\mathcal{Q}$ satisfies properties (34) to (36), and $V=V\circ\varphi$, $M=\varphi^{*}M$, then then all $\zeta(t)$ in the unique LSM $\mathcal{M}\subset T^{*}\mathcal{Q}$ are such that
$\zeta(\mathbb{R})={\sigma}_{2}(\zeta(\mathbb{R}))\,.$ (72)
{pf}
Analogous to the proof of (19), the symmetry $(S_{2},\tau_{2})$ satisfies the conditions of Theorem 5. With
$S_{2}((0,0))=(0,0)\,$, (73)
and eigenspace $E_{0}:=D\oplus M(\bar{x})D$ satisfying
$E_{0}=\frac{\partial S_{2}}{\partial z}E_{0}=\begin{bmatrix}-I&0\\ 0&-I\end{bmatrix}E_{0}\,.$ (74)
Thus, the LSM $\mathcal{M}$ is such that
$\mathcal{M}$ $=\sigma_{2}(\mathcal{M})\,$, (75)
$\zeta(\mathbb{R})$ $=\sigma_{2}\big(\zeta(\mathbb{R})\big)\,.$ (76)
Lemma 25.
Given an evolution $\zeta(t)=(x(t),P(t))$ of (2). If there is a pair of points $(x,P),(\varphi(x),{\varphi^{-1}}^{*}P)\in\zeta(\mathbb{R})$ and denoting $\Delta t=\min\{t>0\,|\,(\varphi(x),{\varphi^{-1}}^{*}P)=\Psi^{t}_{X_{H}}((x,P))\}$, then the evolution is periodic with period $T=2\Delta t$ and
$x(t+T/2)$ $=\varphi\big(x(t)\big)\,$, (77)
$P(t+T/2)$ $={\varphi^{-1}}^{*}P(t)\,.$ (78)
{pf}
Let $\zeta_{1}(t)=(x(t),P(t))$ be a solution to (2), (1). Via (37)
$\zeta_{2}(t)=(\varphi(x)(t),{\varphi^{-1}}^{*}P(t))\,$, (79)
must also be a solution. By hypothesis, the points $\zeta_{1}(0)=(x,P)$ and $\zeta_{2}(0)=(\varphi(x),{\varphi^{-1}}^{*}P)$ lie on the same trajectory, and
$\zeta_{1}(\Delta t)=\zeta_{2}(0)\,.$ (80)
Similarly $\zeta_{2}(\Delta t)=\zeta_{1}(0)$, such that $\zeta_{1}(2\Delta t)=\zeta_{2}(\Delta t)=\zeta_{1}(0)$, i.e., $\zeta(t)$ is periodic and $T=2\Delta t$ is the period. It also follows from (80) that $\zeta_{1}(t+T/2)=\zeta_{2}(t)$, which yields
$x(t+T/2)$ $=\varphi\big(x(t)\big)$ (81)
$P(t+T/2)$ $={\varphi^{-1}}^{*}P(t)\,$, (82)
which concludes the proof.
Lemma 26 (Time to equilibrium).
[…]
{pf}
Since there are two points with $P=0$, Lemma 17 applies and $\zeta(t)$ is periodic with period $T=2\Delta t$. We set $\zeta(0)=(x,0)$ and $\zeta(T/2)=(\varphi(x),0)$. Equation (44) applies, and a particular case is
x(T/4)=x(T/4).x(T/4)=x(-T/4)\,. (83)
Points $(x,0),(\varphi(x),0)\in\zeta(\mathbb{R})$ are of the form required for Lemma 25, so equation (77) holds and a particular case is
$x(T/4)=\varphi\big(x(-T/4)\big)\,.$ (84)
Combining (83) and (84) yields
$x(-T/4)=\varphi\big(x(-T/4)\big)\,$, (85)
i.e., $x(t)$ is guaranteed to move through $\widetilde{x}=x(-T/4)$, which is a fixed point of $\varphi:\mathcal{Q}\rightarrow\mathcal{Q}$ . The statement $(\widetilde{x},P)=\Psi^{T/4}_{X_{H}}((x,0))$ follows directly.
Property 2. of Theorem 12 is a direct consequence of Theorem 24 and Lemma 26.
Lemma 27 (Equilibrium Property).
Given $T>0$. If a $T$-periodic orbit $\zeta(t)$ satisfies
$\zeta(\mathbb{R})={\sigma}_{1}(\zeta(\mathbb{R}))={\sigma}_{2}(\zeta(\mathbb{R}))\,$, (86)
and if the fixed point $\bar{x}$ of ${\sigma}_{1},{\sigma}_{2}$ is the only equilbrium such that $V(\bar{x})<V(x)$ for $(x,0)\in\zeta(\mathbb{R})$. Then
$\bar{x}\in\pi\big(\zeta(\mathbb{R})\big)\,.$ (87)
{pf}
[…]
Appendix B Additional Material
B.1 Canonical charts on $T^{*}\mathcal{Q}$
See this section in the supplementary material [34]. Define coordinates on $\mathcal{Q}$ by a chart $(U,X)$ with $U\subset\mathcal{Q}$ and $X:U\rightarrow\mathbb{R}^{n}$, assigning coordinates to $x\in\mathcal{Q}$ by
$(q_{1},\ldots,q_{n}):=X(x)\,.$ (88)
At the point $x$, this induces a fibre-wise chart $(T_{Q}^{*}\mathcal{Q},{X^{-1}}^{*})$ via the pullback ${X^{-1}}^{*}:T_{Q}^{*}\mathcal{Q}\rightarrow\mathbb{R}^{n}$ that assigns to $P\in T_{Q}^{*}\mathcal{Q}$ the coordinates
$(p_{1},\ldots,p_{n})={X^{-1}}^{*}(P)\,.$ (89)
A canonical chart on $T^{*}U\subset T^{*}\mathcal{Q}$ then assigns to $(x,P)\in T^{*}U$ the coordinates $(q_{1},\ldots,q_{n},p_{1},\ldots,p_{n})$.
B.2 The geometric Hessian of a function
Denote by $\text{d}:C^{\infty}(\mathcal{M})\rightarrow\Gamma(T^{*}\mathcal{M})$ the differential of a function and by $\nabla_{\cdot}\cdot:\mathfrak{X}(\mathcal{M})\times\Gamma(T^{p}_{q}\mathcal{M})\rightarrow\Gamma(T^{p}_{q}\mathcal{M})$ the covariant derivative w.r.t. a given connection on $\mathcal{M}$. The Hessian of $V\in C^{2}(\mathcal{M},\mathbb{R})$ is defined as
$\text{Hess}(V)(X,Y):=\big(\nabla_{X}\text{d}V\big)(Y)\,.$ (90)
In a local chart we denote $\text{d}V=\frac{\partial V}{\partial z^{j}}dz^{j}$, write $X=X^{i}\frac{\partial}{\partial z^{j}}$, $Y=Y^{i}\frac{\partial}{\partial z^{j}}$, and the choice of connection corresponds to a choice of Christoffel Symbols $\Gamma^{k}_{ij}$. Using Einstein summation notation, the Hessian reads
$\text{Hess}(V)(X,Y)=X^{j}\frac{\partial^{2}V}{\partial z^{i}\partial z^{j}}Y^{i}-X^{j}\frac{\partial V}{\partial z^{k}}\Gamma^{k}_{ij}Y^{i}\,.$ (91)
At points $m\in\mathcal{M}$ where $\text{d}V=0$, we have that
$\text{Hess}(V)=\frac{\partial^{2}V}{\partial z^{i}\partial z^{j}}\text{d}z^{i}\text{d}z^{j}\,$, (92)
i.e., $\text{Hess}(V)(X,Y)$ is a $(0,2)$-tensor at critical points of $V$, and is then also independent of the choice of Christoffel symbols $\Gamma^{k}_{ij}$.